# A family of log-correlated Gaussian processes

### v1

@@ Abstract.
A family of log-correlated Gaussian processes indexed by metric spaces are introduced. Given a metric space so that the metric is a measure-definite kernel, a fractional Brownian motion with Hurst index $H\in(0,1/2)$ is known to exist. A log-correlated Gaussian process then arises as the limit of fractional Brownian motions, as the Hurst index $H$ tends to zero, after a non-trivial random shift and then scaled by $H^{-1/2}$. When the index set is $\mathbb{R}_{+}$, for each $H\in(0,1/2)$ the shifted fractional Brownian motion is just a bi-fractional Brownian motion. This choice of shift corresponds to the decomposition of fractional Brownian motion as the sum of a bi-fractional Brownian motion and another smooth Gaussian process introduced by Lei and Nualart, [30], and the decomposition is known to have its counterparts for fractional Brownian motions indexed by metric spaces. The generalized processes when evaluated at a test function have stochastic-integral representations. The introduced processes are also shown to be the scaling limits of certain aggregated models.
1. Introduction and main results
@@
where $g(x,y)$ is a continuous function on ${\mathsf{M}}\times{\mathsf{M}}$. The corresponding Gaussian process is not point-wise defined since $\Gamma(x,x)=\infty$. In the seminal work, Kahane, [26] established a general theory of building the Gaussian multiplicative chaos starting from such kernels. Many developments have followed recently. See the survey of Rhodes and Vargas, [41] and the forthcoming monograph by Berestycki and Powell, [3]. In particular, many efforts have been devoted to specific log-correlated Gaussian processes, among which the most prominent example has been the Gaussian free fields [3, 45]. At the same time, log-correlated Gaussian processes and Gaussian multiplicative chaos have been shown to arise from limit theorems in random matrix theory (e.g. [20, 6, 49]) and mathematical finance (e.g. [2, 36, 13]). Most of the time, ${\mathsf{M}}$ is either a domain of $\mathbb{R}^{n}$, or the circle ${\mathbb{S}}^{1}$. See also the survey [8].
@@
The goal of this paper is to introduce a family of log-correlated Gaussian processes with a covariance function explicitly determined by the metric ${\mathsf{d}}$, on spaces including but not limited to $\mathbb{R}$. For Brownian motion, Paul Lévy already investigated the question of defining Brownian motion on other spaces than $\mathbb{R}$ (see e.g. [31, 35, 16]). Later on, the question of defining fractional Brownian motions on other spaces including Euclidean spaces $\mathbb{R}^{d}$, Euclidean spheres ${\mathbb{S}}^{d}$, and hyperbolic spaces $\mathbb{H}^{d}$ with Hurst index $H\in(0,1/2)$ has been addressed, and these processes indexed by metric spaces can be characterized as centered Gaussian processes $G^{H}$ with ${\mathbb{E}}(G^{H}(x)-G^{H}(y))^{2}={\mathsf{d}}^{2H}(x,y)$ and $G^{H}(o)=0$ for some $o\in{\mathsf{M}}$; see for example [24, 23, 25, 7]. What we introduce here can be viewed as counterparts of fractional Brownian motions with Hurst index $H=0$ (strictly speaking they are better not viewed as the fractional Brownian motion with Hurst index $H=0$, a subtle point we shall elaborate at a later moment).
@@
It has been recently shown that when ${\mathsf{d}}$ is a measure definite kernel, a notion from Robertson and Steger, [42], not only a fractional Brownian motion indexed by ${\mathsf{M}}$ with Hurst index $H\in(0,1/2)$ can be defined, but also it has a stochastic-integral representation with Chentsov-type [44]. Our results are closely related to this unified framework for fractional Brownian motions indexed by metric spaces. The metric ${\mathsf{d}}$ being a measure definite kernel is also our key assumption here.
@@
Recall that if ${\mathsf{d}}$ is of negative type then there exists a Gaussian process $(G(x))_{x\in{\mathsf{M}}}$ such that ${\mathbb{E}}(G(x)-G(y))^{2}={\mathsf{d}}(x,y)$. Every measure definite kernel is of negative type, but the opposite is not true. It has been well-known that if ${\mathsf{d}}$ is of negative type, then so is ${\mathsf{d}}^{\beta}$ for all $\beta\in(0,1)$. A brief review on metrics of negative type, measure definite kernels and their induced Gaussian processes can be found in [25]. It has been recently shown that if ${\mathsf{d}}$ is a measure definite kernel, then so is ${\mathsf{d}}^{\beta}$ for all $\beta\in(0,1)$ [14].
The first main result is the following, of which an extended version is provided in Theorem 4.2.
@@
If Assumption 1.1 is satisfied, then
$\Gamma(x,y)=\log\frac{\mu(A_{x})+\mu(A_{y})}{{\mathsf{d}}(x,y)},x,y\in{\mathsf{M}},$
is the covariance function of a generalized Gaussian process on ${\mathsf{M}}$. More precisely, we have
(1.2) $\Gamma(x,y)=\int_{0}^{\infty}\Gamma_{r}(x,y){\rm d}r$
where for each $r>0$, $\Gamma_{r}:{\mathsf{M}}\times{\mathsf{M}}\to\mathbb{R}$ is a positive definite function.
@@
Note that $\Gamma$ in (1.2) is a $\sigma$-positive kernel in the sense of Kahane, [26], a property that is not easy to check in practice. The $\sigma$-positivity is a consequence of the metric ${\mathsf{d}}$ being a measure definite kernel (see (4.3); the expression of $\Gamma_{r}$ depends on $(A_{x})_{x\in{\mathsf{M}}}$ and $\mu$). Thanks to this property, the existence of the Gaussian multiplicative chaos determined by $\Gamma$ follows immediately from Kahane’s framework. We leave the investigations of so-obtained Gaussian multiplicative chaos for future studies.
We shall provide an explicit expression for $\Gamma_{r}$ in Section 4, Theorem 4.2. In fact, in examples of our interest, the kernel of the covariance function shall be uniquely determined by the metric. For example, if there is a point $o\in{\mathsf{M}}$ such that $\mu(A_{o})=0$, then the kernel simplifies to
(1.3) $\Gamma(x,y)=\log\frac{{\mathsf{d}}(x,o)+{\mathsf{d}}(y,o)}{{\mathsf{d}}(x,y)}.$
@@
In the case ${\mathsf{M}}=\mathbb{R}$, we can take $A_{x}=[0,x]$ if $x\geq 0$ and $A_{x}=[x,0]$, and $\mu$ is the Lebesgue measure. Then, we have
(1.4) $\Gamma(x,y)=\log\frac{|x|+|y|}{|x-y|},x,y\in\mathbb{R}.$
Obviously, $\Gamma(x,y)=0$ if $x<0<y$, which means that the process restricted to $(-\infty,0)$ and $(0,\infty)$ respectively are now independent. Therefore, later on for one-dimensional example we shall focus on the generalized process restricted to $\mathbb{R}_{+}$.
@@
For general ${\mathsf{M}}=\mathbb{R}^{n}$, it is well-known that the Euclidean metric ${\mathsf{d}}$ is a measure-definite kernel. Indeed, one can take $E:=\{(s,r):s\in{\mathbb{S}}^{n},r\in(0,\infty)\}$ representing the space of all hyperplanes of $\mathbb{R}^{n}$ not including the origin $A_{x}=\{(s,r):s\in{\mathbb{S}}^{n},r\in(0,\langle s,x\rangle)\}$ the collection of all hyperplanes separating 00 and $x$ in $\mathbb{R}^{n}$, and $\mu({\rm d}s{\rm d}r)={\rm d}s{\rm d}r$ (${\rm d}s$ is the Lebesgue measure on ${\mathbb{S}}^{n}$, ${\rm d}r$ is the Lebesgue measure on $\mathbb{R}^{n}$, such that (1.1) holds; see [44, Chapter 8.3]. So the covariance kernel is
$\Gamma(x,y)=\log\frac{{\left\|x\right\|}+{\left\|y\right\|}}{{\left\|x-y\right\|}},\quad x,y\in\mathbb{R}^{n}.$
@@
In the case ${\mathsf{M}}={\mathbb{S}}^{n}$ and ${\mathsf{d}}$ is the geodesic, we can take $H_{x}$ as the hemisphere of ${\mathbb{S}}^{n}$ centered as $x\in{\mathbb{S}}^{n}$, $o\in{\mathbb{S}}^{n}$ arbitrarily chosen and then fixed as an origin, $A_{x}:=H_{x}\Delta H_{o},x\in{\mathbb{S}}^{n}$ and $\mu({\rm d}x)=\pi{\rm d}x/w_{n}$ where $w_{n}$ is the surface area of ${\mathbb{S}}^{n}$. In this way, (1.1) and (1.3) hold.
Note that the above choice is well-known when studying spherical fractional Brownian motions. Note that the spherical fractional Brownian motions and the corresponding log-correlated Gaussian processes are not rotation-invariant. For a rotation-invariant model, one can consider simply $A_{x}=H_{x}$. This time, we still have ${\mathsf{d}}(x,y)=\mu(H_{x}\Delta H_{y})$ (since $(H_{x}\Delta H_{o})\Delta(H_{y}\Delta H_{o})=H_{x}\Delta H_{y}$). But now we have $\mu(H_{x})=\pi/2$ (compared with $\mu(A_{x})={\mathsf{d}}(o,x)$) and hence the corresponding generalized Gaussian process has covariance kernel
$\Gamma(x,y)=\log\frac{\pi}{{\mathsf{d}}(x,y)}.$
@@
For the rest of the introduction and the majority of the paper (except Section 4), we shall focus on Example 1.1, the log-correlated Gaussian process on $\mathbb{R}_{+}=[0,\infty)$ with covariance function (1.4) (restricted to $\mathbb{R}_{+}\times\mathbb{R}_{+}$). This is the first example we discovered for the general framework leading to Theorem 1.4, and it relates to several recent developments regarding fractional Brownian motions. Focusing on this example shall help illustrate all the key ideas behind the general framework under Assumption 1.1, and the extension to the general framework should become straightforward after Example 1.1 is thoroughly examined.
1.1. From fractional Brownian motions to logarithmic-correlated Gaussian processes
We start by explaining the relation between our introduced processes and fractional Brownian motions. Our studies were inspired by Neuman and Rosenbaum, [36] who, to the best of our knowledge, first considered the question of limit process of fractional Brownian motion as $H\downarrow 0$. See Remark 1.7 below for more details on their approach. The connection of our process $G$ to fractional Brownian motions is different from theirs. The revealed connection here is in a sense more generic as it can be extended in a unified approach under Assumption 1.1 to relate a log-correlated Gaussian processes to fractional Brownian motions indexed by other metric spaces.
We first recall some facts on fractional Brownian motions. We let $({\mathbb{B}}^{H}_{t})_{t\in[0,\infty)}$ denote a fractional Brownian motion on $[0,\infty)$ with Hurst index $H\in(0,1)$. It is known that this is the centered Gaussian process with covariance function
${\rm{Cov}}\left({\mathbb{B}}_{s}^{H},{\mathbb{B}}_{t}^{H}\right)=\frac{1}{2}\left(s^{2H}+t^{2H}-|s-t|^{2H}\right),\quad s,t\geq 0.$
One may compute directly the limit of the covariance function, and it is obvious that a non-trivial shift of ${\mathbb{B}}^{H}$ is necessary to obtain a non-degenerate limit. Our solution is related to the decomposition of fractional Brownian motion into the sum of a bi-fractional Brownian motion and a smooth Gaussian process due to Lei and Nualart, [30]. Namely, we can write
(1.5) $\left({\mathbb{B}}^{H}_{t}\right)_{t\in[0,\infty)}\stackrel{{\scriptstyle d}}{{=}}\left(G^{H}_{t}+W^{H}_{t}\right)_{t\in[0,\infty)},$
where the two processes $G^{H}$ and $W^{H}$ are independent centered Gaussian ones with covariance functions
${\rm{Cov}}\left(G^{H}_{s},G^{H}_{t}\right)$ $=\frac{1}{2}\left((s+t)^{2H}-|t-s|^{2H}\right),$
${\rm{Cov}}\left(W^{H}_{s},W^{H}_{t}\right)$ $=\frac{1}{2}\left(s^{2H}+t^{2H}-(s+t)^{2H}\right).$
The process $G^{H}$ is a special case of so-called bi-fractional Brownian motions (see e.g. [10, Remarks 2.4 and 2.5] and [48, 32] for more discussions), and $W^{H}$ is a smooth Gaussian process. (In fact, the processes ${\mathbb{B}}^{H},G^{H},W^{H}$ can be extended to processes indexed by $\mathbb{R}$, and the decomposition remains to hold. However, the restrictions of process $G^{H}$ to $(-\infty,0)$ and $(0,\infty)$ respectively are again independent, and hence it suffices to focus on $[0,\infty)$.)
Now, it is clear that
(1.6) $\lim_{H\downarrow 0}\frac{{\rm{Cov}}(G^{H}(s),G^{H}(t))}{H}=\Gamma(s,t)=\log\frac{s+t}{|s-t|},\quad s,t\geq 0,s\neq t.$
(And there is no corresponding convergence for the covariance function of ${\mathbb{B}}^{H}$ and $W^{H}$.) Accordingly, we have formally $H^{-1/2}G^{H}\Rightarrow G$ as random generalized functions.
The process $G^{H}$ can be viewed as ${\mathbb{B}}^{H}$ shifted by $W^{H}$. It does not make sense immediately to write ${\mathbb{B}}^{H}_{t}-W^{H}_{t}$ as the dependence structure cannot be specified from (1.5): one can do so after introducing a stochastic-integral representation of the decomposition in Section 2.
The limit covariance function in (1.6) can be taken as the definition of the log-correlated process. But this is not the end of the story. At the heart of aforementioned recent developments of fractional Brownian motions is a new stochastic-integral representation, which reflects the aggregation nature of the processes. It turned out that $G(f)$ also has a stochastic-integral representation $f$ from a suitable family ${\mathcal{F}}_{0}$ of continuous functions (see Lemma 2.1). (Note that this is not the same as saying that one can write $G(f)=\int_{0}^{\infty}\widetilde{G}(t)f(t){\rm d}t$ for some stochastic process $\widetilde{G}(t)$, and in fact such an expression does not hold.) A few related remarks shall be made immediately in Section 2. In particular, as an immediate consequence we have a natural extension of $G$ to a generalized stable process; see Lemma 2.3.
@@ 1.1. From fractional Brownian motions to logarithmic-correlated Gaussian processes
In fact, the generalized Gaussian process $G$ can be extended to a generalized infinitely-divisible process [43] by replacing in the stochastic-integral representation the Gaussian random measures by an infinitely-divisible one (see [2, 40] for a closely related example in a similar spirit). We do not pursue the full generality in this direction here. Instead, as will be shown later the extension to stable ones is almost free, and also the limit theorems can be also extended accordingly. It is also worth pointing out that for fractional Brownian motions indexed by metric spaces and their stable counterparts, they all have stationary increments in the sense that the law of the increment of the process $Z(x)-Z(y)$ depends only on ${\mathsf{d}}(x,y)$ (see [25, 14]). It is not clear whether a similar invariance property holds for the generalized process $G$.
For the rest of the introduction, we just mention that $G$ is a log-correlated Gaussian process with covariance function (1.4), in the sense that $\{G(f)\}_{f\in{\mathcal{F}}_{0}}$ is a family of centered Gaussian random variables represented via stochastic integrals and satisfying
(1.7) ${\rm{Cov}}(G(f),G(g))=\int_{0}^{\infty}\int_{0}^{\infty}f(s)g(t)\log\frac{s+t}{|s-t|}{\rm d}s{\rm d}t,\quad\mbox{ for all }f,g\in{\mathcal{F}}_{0},$
with
${\mathcal{F}}_{0}:=\left\{f\in L^{\infty}([0,\infty)):\int_{0}^{\infty}s^{\delta}|f(s)|{\rm d}s<\infty\mbox{ for some $\delta>0$}\right\}.$
In view of the connection to the decomposition of fractional Brownian motions (1.6), the introduced process $G$ is better viewed as a log-correlated bi-fractional Brownian motion. This name however is also a little misleading as bi-fractional Brownian motions are indexed by two parameters and $G^{H}$ is with a very specific choice of the parameters. Instead, the most crucial feature in all the developments here is certain ‘odd-occupancy event’ in the stochastic-integral representation in (2.2) (see Lemma 2.1), and it is not clear whether other bi-fractional Brownian motions have a similar representation.
In fact, the discussions on fractional Brownian motions above are closely related to earlier investigations starting from [10] on the so-called odd-occupancy processes of the Karlin model [28, 17], of which the scaling limits lead to the decomposition of fractional Brownian motions in (1.5). For this reason we propose to name the process $G$ as the log-correlated Karlin (Gaussian) process. We shall briefly discuss the connection of our results with Karlin model in Remark 1.9.
1.2. A central limit theorem
Another feature of the introduced process $G$ is that it arises from a simple aggregation model. The aggregation model here is a variation of the one proposed by Enriquez, [11] when investigating fractional Brownian motions with $H\in(0,1/2)$. Connections between aggregated models and fractional Gaussian processes and random fields with long-range dependence [38, 43] are well-known, see for example [4, 5, 27, 34] (in particular, [4, 5] established limit theorems for generalized processes).
We follow the notations in [46], where the Enriquez model was revisited and extended. The model is an aggregated model of an increasing number, denoted by $m_{n}$, of i.i.d. layers indexed by $i\in\mathbb{N}$, and each layer consists of a sequence of stationary random variables denoted by $(\mathcal{X}_{i,j})_{j\in\mathbb{N}}$ that we now introduce. First, let $\mathcal{X}$ be a standard Gaussian random variable, $q$ a uniform random variable taking values from $(0,1)$, and $\{\eta_{j}^{(q)}\}_{j\in\mathbb{N}}$ conditionally i.i.d. Bernoulli random variables with parameter $q$. Assume that $\mathcal{X}$ is independent from $q$ and $\{\eta_{j}^{(q)}\}_{j\in\mathbb{N}}$. Let $(\mathcal{X}_{i},q_{i},\{\eta_{i,j}^{(q_{i})}\}_{j\in\mathbb{N}})_{i\in\mathbb{N}}$ denote i.i.d. copies of $(\mathcal{X},q,\{\eta_{j}^{(q)}\}_{j\in\mathbb{N}})$, and set
$\tau_{i,j}^{(q_{i})}:=\sum_{k=1}^{j}\eta_{i,k}^{(q_{i})},j\in\mathbb{N}.$
Then, the $j$-th element of $i$-th layer is defined as
In order to see the proposed model exhibits log-correlated dependence structure, a quick calculation yields
(1.8) $\lim_{n\to\infty}{\rm{Cov}}\left(\widehat{S}_{n}(s),\widehat{S}_{n}(t)\right)=\frac{1}{4}\log\left(\frac{s+t}{s-t}\right),0<s<t.$
Indeed,
$\to\frac{1}{4}\int_{0}^{\infty}\frac{1}{r}\left(e^{-2(t-s)r}-e^{-2(t+s)r}\right){\rm d}r=\frac{1}{4}\log\left(\frac{t+s}{t-s}\right).$
The convergence in (1.8) suggests that there is no convergence of point-wise defined stochastic processes, and instead one should search for convergence in the sense of generalized processes. For this purpose, introduce
(1.9) $G_{n}(f):=2\sum_{j=1}^{\infty}\widehat{S}_{n,j}\int_{(j-1)/n}^{j/n}f(x){\rm d}x,\quad f\in{\mathcal{F}}_{0}.$
(Note that we do not assume $f$ to be continuous for $f\in{\mathcal{F}}_{0}$. If continuity is assumed, then one can replace $\int_{(j-1)/n}^{j/n}f(x){\rm d}x$ by $f(j/n)/n$ and establish the same limit theorem below.)
The second main result is the following limit theorem.
Theorem 1.4.
Assume $\lim_{n\to\infty}m_{n}/n=\infty$. Then, for all $f\in{\mathcal{F}}_{0}$,
[…]

### v3

@@ Abstract.
A family of log-correlated Gaussian processes indexed by metric spaces is introduced, when the metric is conditionally negative definite. These processes arise as the limit of bi-fractional Brownian motions indexed by $(H,K)$ scaled by $K^{-1/2}$ as $K\downarrow 0$ with $H\in(0,1/2]$ fixed. When the metric is in addition a measure definite kernel, stochastic-integral representations of the generalized processes when evaluated at a test function are provided. The introduced processes are also shown to be the scaling limits of certain aggregated models.
1. Overview
@@
where $g(x,y)$ is a continuous function on ${\mathsf{M}}\times{\mathsf{M}}$. The corresponding Gaussian process is not point-wise defined since $\Gamma(x,x)=\infty$. In the seminal work, Kahane, [26] established a general theory of building the Gaussian multiplicative chaos starting from such kernels. Many developments have followed recently. See the survey of Rhodes and Vargas, [42] and the forthcoming monograph by Berestycki and Powell, [3]. In particular, many efforts have been devoted to specific log-correlated Gaussian processes, among which the most prominent example has been the Gaussian free fields [3, 46]. At the same time, log-correlated Gaussian processes and Gaussian multiplicative chaos have been shown to arise from limit theorems in random matrix theory [20, 6, 50] and mathematical finance [2, 37, 13]. Most of the time, $\mathsf{M}$} is either a domain of $\mathbb{R}^{n}$, or the circle ${\mathbb{S}}^{1}$. See also the survey by Duplantier et al., [8].
@@
The goal of this paper is to introduce a family of log-correlated Gaussian processes indexed by metric spaces with a covariance function explicitly determined by the metric $\mathsf{d}$} and a parameter $H\in(0,1/2]$. For Brownian motion, Paul Lévy already investigated the question of defining Brownian motion indexed by other spaces than $\mathbb{R}$ [31, 36, 16]. Later on, the question of defining fractional Brownian motions indexed by other spaces including Euclidean spaces $\mathbb{R}^{d}$, Euclidean spheres ${\mathbb{S}}^{d}$, and hyperbolic spaces $\mathbb{H}^{d}$ with Hurst index $H\in(0,1/2)$ has been addressed, and these processes indexed by metric spaces can be characterized as centered Gaussian processes $G^{H}$ with ${\mathbb{E}}(G^{H}(x)-G^{H}(y))^{2}={\mathsf{d}}^{2H}(x,y)$ and $G^{H}(o)=0$ for some $o\in{\mathsf{M}}$ [24, 23, 25, 7]. A closely related question is the existence of the so-called bi-fractional Brownian motions indexed by metric spaces. A bi-fractional Brownian motion is a Gaussian process with two parameters that we let $H,K$ denote throughout. We shall see that for a large family of examples with $H\in(0,1/2]$, a log-correlated Gaussian process arises as the simple limit of bi-fractional Brownian motions with parameters $(H,K)$ normalized by $K^{-1/2}$ as $K\downarrow 0$. This, actually, follows from a simple calculation that we now explain.
1.1. From the bi-fractional Brownian motion to a log-correlated Gaussian process
We let $G^{H,K}=(G^{H,K}(x))_{x\in{\mathsf{M}}}$ denote the bi-fractional Brownian motion indexed by a metric space $\mathsf{M}$}, and $H>0,K>0$ are two parameters that we shall discuss in a moment. This is a centered Gaussian process with covariance function
(1.1) ${\rm{Cov}}\left(G^{H,K}(x),G^{H,K}(y)\right)=\frac{1}{2^{K}}\left(\left({\mathsf{d}}^{2H}(o,x)+{\mathsf{d}}^{2H}(o,y)\right)^{K}-{\mathsf{d}}^{2HK}(x,y)\right),x,y\in{\mathsf{M}},$
where $o\in{\mathsf{M}}$ is fixed and plays the role of the origin. Bi-fractional Brownian motion was first introduced by Houdré and Villa, [19] with $\mathsf{M}$$\mathbb{R}$ (and in this case $G^{H,K}$ is well-defined with $H\in(0,1],K\in(0,1]$ and $o=0$). In particular, $G^{H,1}$ is just a fractional Brownian motion with Hurst index $H\in(0,1]$ ($H=1$ is a degenerate case). The exact range of legitimate $(H,K)$ for the existence of $G^{H,K}$ remains an open question [49] even when $\mathsf{M}$$\mathbb{R}$ and $\mathsf{d}$} is the Euclidean metric; see also Lifshits and Volkova, [32].
Since then, bi-fractional Brownian motions indexed by various metric spaces have been studied; see Ma, [33]. In particular, when setting $K=1$ one always recover a fractional Brownian motion indexed by the same metric space with Hurst index $H$. It is important to keep in mind that the legitimate choice of $(H,K)$ depends on $({\mathsf{M}},{\mathsf{d}})$. As a general setup we assume that $\mathsf{d}$} is conditionally negative definite (a.k.a. of negative type), and in this case it is easy to check that the function on the right-hand side of (1.1) is positive definite and hence $G^{H,K}$ is well-defined with
$H\in(0,1/2]\quad\mbox{ and }\quad K\in(0,1].$
See Ma, [33, Theorem 1.1, (vii)]. In particular, $G^{H,1}$ is a fractional Brownian motion with Hurst index $H$ and $G^{1/2,1}$ is a Brownian motion.
Then, with $H$ fixed we have
(1.2) $\lim_{K\downarrow 0}\frac{1}{K}{\rm{Cov}}\left(G^{H,K}(x,y)\right)=\log\frac{{\mathsf{d}}^{2H}(o,x)+{\mathsf{d}}^{2H}(o,y)}{{\mathsf{d}}^{2H}(x,y)}=:\Gamma^{(H)}(x,y),x,y\in{\mathsf{M}}.$
This simple calculation deserves to be stated in a proposition.
Proposition 1.1.
If $({\mathsf{M}},{\mathsf{d}})$ is a metric space and $\mathsf{d}$} is of negative type, then for all $H\in(0,1/2]$, $\Gamma^{(H)}$ in (1.2) is a covariance kernel for a generalized Gaussian process indexed by $\mathsf{M}$}.
We let $G^{H}$ denote the so-introduced Gaussian process, and name it as the log-correlated bi-fractional Brownian motion with parameter $H\in(0,1/2]$.
Our first main contribution is to provide a stochastic-integral representation of $G^{H}$. Strictly speaking, we represent $G^{H}(f)$ (when evaluated at a suitable test function $f:{\mathsf{M}}\to\mathbb{R}$) in the form of $G^{H}(f)=\int_{\mathsf{M}}f(x)M({\rm d}x)$ where $M$ is a suitable Gaussian random measure. For this purpose, we shall impose a further condition on $\mathsf{d}$} being of negative type.
@@ 1.1. From the bi-fractional Brownian motion to a log-correlated Gaussian process
It has been recently shown that when $\mathsf{d}$} is a measure definite kernel, a notion from Robertson and Steger, [43], not only a fractional Brownian motion indexed by $\mathsf{M}$} with Hurst index $H\in(0,1/2)$ can be defined, but also it has a stochastic-integral representation of Chentsov-type [45]. Our results are closely related to this unified framework for fractional Brownian motions indexed by metric spaces [14]. The metric $\mathsf{d}$} being a measure definite kernel is also our key assumption here.
@@
Recall that if $\mathsf{d}$} is of negative type then there exists a Gaussian process $(G(x))_{x\in{\mathsf{M}}}$ such that ${\mathbb{E}}(G(x)-G(y))^{2}={\mathsf{d}}(x,y)$. Every measure definite kernel is of negative type, but the opposite is not true. It has been well-known that if $\mathsf{d}$} is of negative type, then so is ${\mathsf{d}}^{\beta}$ for all $\beta\in(0,1)$. A brief review on metrics of negative type, measure definite kernels and their induced Gaussian processes can be found in Istas, [25]. It has been recently shown that if $\mathsf{d}$} is a measure definite kernel, then so is ${\mathsf{d}}^{\beta}$ for all $\beta\in(0,1)$ [14].
The first main result is the following.
@@
If Assumption 1.2 is satisfied, then for $H\in(0,1/2]$,
$\Gamma^{(H)}(x,y)=\log\frac{\mu^{2H}(A_{x})+\mu^{2H}(A_{y})}{{\mathsf{d}}^{2H}(x,y)},x,y\in{\mathsf{M}},\beta\in(0,1],$
is the covariance kernel of a generalized Gaussian process on $\mathsf{M}$}. More precisely, we have
(1.4) $\Gamma^{(H)}(x,y)=\int_{0}^{\infty}\Gamma_{r}^{(H)}(x,y){\rm d}r$
where for each $r>0$, $\Gamma_{r}^{(H)}:{\mathsf{M}}\times{\mathsf{M}}\to\mathbb{R}$ is a positive definite function.
We shall provide an explicit expression for $\Gamma_{r}^{(H)}$, which is essentially based on a stochastic-integral representation of $G(f)$, in Section 4, Theorem 4.2.
@@
Note that $\Gamma^{(H)}$ in (1.4) is a $\sigma$-positive kernel in the sense of Kahane, [26], a property that is not easy to check in practice. The $\sigma$-positivity is a consequence of the metric $\mathsf{d}$} being a measure definite kernel (see (4.3)). Thanks to this property, the existence of the Gaussian multiplicative chaos determined by $\Gamma$ follows immediately from Kahane’s framework. We leave the investigations of so-obtained Gaussian multiplicative chaos for future studies.
Below we recall a few examples satisfying Assumption 1.2. If there is a point $o\in{\mathsf{M}}$ such that $\mu(A_{o})=0$, then the kernel simplifies to $\Gamma^{(H)}$ in (1.2). We also provide an example that satisfies Assumption 1.2 but cannot be obtained as the limit of bi-fractional Brownian motions in Example 1.3 below.
@@
In the case $\mathsf{M}$$\mathbb{R}$, we take $A_{x}=[0,x]$ if $x\geq 0$ and (x,0](-x,0] if $x<0$, and $\mu$ as the Lebesgue measure. Then, we have
(1.5) $\Gamma^{(H)}(x,y)=\log\frac{|x|^{2H}+|y|^{2H}}{|x-y|^{2H}},x,y\in\mathbb{R}.$
Obviously, when $H=1/2$, $\Gamma(x,y)=0$ if $x<0<y$, which means that the restrictions of the process to $(-\infty,0)$ and $(0,\infty)$ respectively are independent. In Sections 2 and 3 we shall focus on this example with $H=1/2$ and restricted to $0,\infty)$.
@@
For ${\mathsf{M}}=\mathbb{R}^{n}$, it is well-known that the Euclidean metric $\mathsf{d}$} is a measure-definite kernel. Indeed, one can take $E:=\{(s,r):s\in{\mathbb{S}}^{n},r\in(0,\infty)\}$ representing the space of all hyperplanes of $\mathbb{R}^{n}$ not including the origin, $A_{x}=\{(s,r):s\in{\mathbb{S}}^{n},0<r<\langle s,x\rangle\}$ the collection of all hyperplanes separating 0 and $x$ in $\mathbb{R}^{n}$, and $\mu({\rm d}s{\rm d}r)={\rm d}s{\rm d}r$ (${\rm d}s$ is the Lebesgue measure on ${\mathbb{S}}^{n}$, ${\rm d}r$ is the Lebesgue measure on $\mathbb{R}^{1}$, such that (1.3) holds. In fact, this representation of the measure definite kernel is at the heart of earlier Chentsov representation for fractional Lévy Brownian fields [45, Chapter 8.3]. So the covariance kernel is
$\Gamma^{(H)}(x,y)=\log\frac{{\left\|x\right\|}^{2H}+{\left\|y\right\|}^{2H}}{{\left\|x-y\right\|}^{2H}},\quad x,y\in\mathbb{R}^{n}.$
@@
In the case ${\mathsf{M}}={\mathbb{S}}^{n}$ and $\mathsf{d}$} is the geodesic, we can take ${\mathcal{H}}_{x}$ as the hemisphere of ${\mathbb{S}}^{n}$ centered as $x\in{\mathbb{S}}^{n}$, $o\in{\mathbb{S}}^{n}$ a fixed point, $A_{x}:={\mathcal{H}}_{x}\Delta{\mathcal{H}}_{o},x\in{\mathbb{S}}^{n}$ and $\mu({\rm d}x)=\pi{\rm d}x/w_{n}$ where $w_{n}$ is the surface area of ${\mathbb{S}}^{n}$. In this way, (1.3) and (1.2) follow.
The above choice is well-known when studying the spherical fractional Brownian motion. Note that the spherical fractional Brownian motion and the corresponding log-correlated Gaussian process are not rotation-invariant. For a rotation-invariant model, one can consider simply $A_{x}={\mathcal{H}}_{x}$. This time, we still have ${\mathsf{d}}(x,y)=\mu({\mathcal{H}}_{x}\Delta{\mathcal{H}}_{y})$ (since $({\mathcal{H}}_{x}\Delta{\mathcal{H}}_{o})\Delta({\mathcal{H}}_{y}\Delta{\mathcal{H}}_{o})={\mathcal{H}}_{x}\Delta{\mathcal{H}}_{y}$). But now we have $\mu({\mathcal{H}}_{x})=\pi/2$ (compared with $\mu(A_{x})={\mathsf{d}}(o,x)$) and hence the corresponding generalized Gaussian process has covariance kernel
$\Gamma^{(H)}(x,y)=\log\frac{2(\pi/2)^{2H}}{{\mathsf{d}}^{2H}(x,y)}.$
@@
The fact that $G^{H}(f)$ has a stochastic-integral representation with $f$ from a suitable family of test functions yields immediately that the generalized Gaussian process $G$ can be extended to generalized infinitely-divisible processes [44] by replacing in the stochastic-integral representation the Gaussian random measure by infinitely-divisible ones (see [2, 41] for a closely related example in a similar spirit). We do not pursue the full generality in this direction here. Instead, we explain that the extension to stable ones is almost free. Last but not least, as the second main contribution we provide limit theorems for the introduced log-correlated bi-fractional Brownian motions.
The paper is organized as follows. A unified treatment for general $({\mathsf{M}},{\mathsf{d}})$ satisfying Assumption 1.2 and with $H\in(0,1/2]$ is provided in Section 4. However, the presentation might be a little too abstract for the first time reading. To help illustrate, we shall focus on Example 1.1, the log-correlated bi-fractional Brownian motion on $\mathbb{R}_{+}=[0,\infty)$ with covariance function (1.5) and $H=1/2$. For this case, in Section 2 we focus on stochastic-integral representations, and in Section 3 as the second main contribution we provide a central limit theorem based on a variation of a model by Enriquez, [11] when studying the fractional Brownian motion. This is also the case closely related to a few very recent developments in the literature; see Remarks 1.5 and 1.6. In Section 4, after presenting the results in the general framework, we also provide a central limit theorem for $G$ based on another model similar to the one in Fu and Wang, [14].
1.2. Related results in the literature
We conclude the introduction with a few remarks.